If g(x)=x^2 and f(x)= 2x+1, then g(f(x))
Answer by stanbon(75887)

If g(x)=x^2 and f(x)= 2x+1, then g(f(x))
---------
Evaluate "g" at f(x).
-----
g[f(x)] = g[x^2] = 2x^2+1
=============================
Cheers,
Stan H.